The area of a sector of a circle with the radius 7cm is 15 square cm. Find the measure of the central angle of the sector in degrees.
Answer by Theo(13342)

the area of a sector is equal to x / 360 * pi * r^2.

your area is 15 square cm.

your radius is equal to 7 cm.

the formula becomes 15 square cm = x / 360 * pi * 7^2

solve for x to get x = 15 * 360 / pi / 7^2 which becomes x = 35.07904868 degrees.

to see if that's correct, go back to the original formula and replace x with 35.07904868 and solve for the area to see is you get 15 square cm.

the formula is:

A = x/360 * pi * r^2

x = 35.07904868
r = 7
r^2 = 49

formula becomes A = (35.07904868 / 360) * pi * 49

solve for A to get A = 15.

solution is confirmed as good.

angle = 35.07904868 degrees.